OpenStudy (anonymous):

The hypotenuse of a 30°-60°-90° triangle measures 32 in. What is the length of the shorter leg?

OpenStudy (anonymous):

 now according to the angles... the shortest side is "b" and we can find it by using sin of 30 degrees \[\LARGE \sin30^{\circ}=\frac{b}{c}\] since.. \[\LARGE \sin30^{\circ}=\frac12\] we have: \[\LARGE \frac12=\frac{b}{32}\] \[\LARGE 2b=32 \Longrightarrow b=\frac{32}{2} \longrightarrow b=16\]